# Temporal-Spectral-Spatial Unified Remote Sensing Dense Prediction

### v1

@@ IV-A Datasets
We offer a brief description of the experimental building and LULC scene datasets in Table I.
1) Building scene datasets: The WHU Building dataset [37] is divided into two main components: one containing satellite imagery and another composed of aerial photos. In our study, we utilize the aerial photo subset, which consists of 8,189 images. These images are split into 4,736 for training, 1,036 for validation, and 2,416 for testing, each with a spatial resolution of 0.3 meters. In total, this subset represents over 22,000 buildings covering an area in excess of 450 square kilometers. Our experiments were conducted using the original partitioning scheme and image dimensions (512×512) as specified by the WHU dataset.
The WHU-CD dataset [37] includes bitemporal very high-resolution (VHR) aerial images taken in 2012 and 2016, which clearly highlight major changes in building structures. The dataset is partitioned into non-overlapping patches of 1024×1024 pixels. These patches are further allocated into training, validation, and test sets following a 7:1:2 ratio.
@@
2) LULC scene datasets: The LoveDA dataset [39] consists of 5,987 high-resolution optical remote sensing images (with a ground sampling distance of 0.3 m) each sized at 1024×1024 pixels. It covers seven land cover classes: building, road, water, barren, forest, agriculture, and background. The dataset is divided into 2,522 training images, 1,669 images for validation, and 1,796 images for testing, all drawn from two distinct scenes—urban and rural—from three Chinese cities: Nanjing, Changzhou, and Wuhan. The dataset poses considerable challenges due to the presence of multiscale objects, complex backgrounds, and uneven class distribution.
@@ IV-B Baseline
To evaluate the effectiveness of the proposed TSSUN, we conducted comparative experiments with various benchmark methods on the building and LULC scene datasets. The benchmark methods tested on the same dataset are based on the same splitting of the dataset and use the same data.
On the four building scene datasets, the compared CNN-based models include FCN [13], SegNet [41], U-Net [14], PSPNet [42], HRNet [43], MA-FCN [44], Deeplabv3+ [16], ResUNet [45], MAPNet [46], D-LinkNet [47], SIINet [48], FC-EF [20], FC-Siam-Diff [20], FC-Siam-Conc [20], STANet [25], DTCDSCN [49], SNUNet [50], CDNet [51], DDCNN [52], DASNet [53], DSIFN [54], HANet [55], USSFCNet [56], and SEIFNet [57], the compared transformer-based models include Segformer [58], ChangeFormer [59] and A2Net [60], and the CNN-transformer hybrid models include BDTNet [61], TransUNet [62], CMTFNet [63], BIT [64], MTCNet [65], MSCANet [66], AMTNet-50 [67], Contrast-COUD [38] and TS-COUD [38].
@@ IV-C Implementation Details
2) Training and Inference: The TSSUN model was implemented using PyTorch [83] and executed on a single RTX A100 GPU (80G). Due to the heterogeneous image resolutions across the datasets, the batch size was set to 16 for the four building scene datasets and 4 for the two LULC scene datasets. Our optimization strategy combined binary cross-entropy loss with Dice coefficient loss, facilitating a balanced performance optimization. The AdamW optimizer [84] was initialized with a learning rate of 0.0001 and a weight decay of 0.001. A learning rate scheduler was employed to reduce the learning rate by a factor of 0.1 if no increase in the mean F1-score was observed on the aggregate validation set for 5 consecutive epochs. The training process spanned 100 epochs, ensuring robust convergence, and the best performing checkpoints—corresponding to the maximum mean F1-scores achieved—were retained for the testing phase. Furthermore, in order to ensure comparability with existing methodologies, all models were initialized using the default PyTorch settings across all datasets.
@@ IV-D Ablation Study
To ascertain the efficacy of the proposed Temporal-Spectral-Spatial Unified Strategy (TSSUS) and Local-Global Window Attention (LGWA), ablation studies were performed on the TSCD dataset. As comparative baselines, TSSUS was replaced with a strategy involving the direct unification of temporal, spectral, and spatial dimensions at both the input and output stages ( referred to as ”direct unification”). Concurrently, LGWA was substituted with standard global attention. This experimental design serves to highlight the distinct advantages of TSSUS in its independent extraction of temporal features and the capability of LGWA in the simultaneous capture of both local and global contextual information.
TABLE II: Accuracy Comparison on the WHU Dataset. The Best Values Are Highlighted in Bold.
Strategy Attention F1 IoU OA
TSSUS LGWA 66.48 49.79 98.05
Direct Unification LGWA 63.22 46.22 94.72
TSSUS Global Attention 65.91 49.15 97.21
Direct Unification Global Attention 62.87 45.85 94.38
Table 2 presents the results of these ablation studies, indicating that both TSSUS and LGWA yield superior performance compared to their respective baseline alternatives. Specifically, TSSUS distinguishes itself from the direct unification approach. It begins by unifying the spectral and spatial dimensions at the input stage. Then, within the encoder, it independently extracts features for each temporal instance (i.e., each remote sensing image). Subsequently, these temporally distinct features are integrated via feature-level fusion at the interface between the encoder and decoder. This methodology—processing temporal information independently before fusion—effectively minimizes interference from irrelevant pixel information within the remote sensing imagery. Furthermore, TSSUS dynamically adjusts the temporal length of feature fusion to align with the specific requirements of the target dense prediction task. Following this, it optimizes the output for each time step independently within the decoder, thereby significantly enhancing the overall model efficacy.
@@ IV-E Overall Comparison
IV-E1 Building
The efficacy of the proposed Temporal-Spectral-Spatial Unified Network was rigorously evaluated through extensive experiments on four benchmark remote sensing datasets: the WHU dataset for single-temporal semantic segmentation, the WHU-CD dataset for bi-temporal semantic change detection, the LEVIR-CD dataset for bi-temporal binary change detection, and the TSCD dataset for multi-temporal binary change detection. Our method was compared against several state-of-the-art approaches, with quantitative results summarized in Table III, IV, V and VI.
On the WHU dataset, characterized by its complex building footprints and significant variations in object scale, our proposed TSSUN achieves state-of-the-art performance. As detailed in Table III, TSSUN obtains the highest IoU of 91.00% and an F1-score of 95.29%. This superior performance can be attributed to TSSUN’s Temporal-Spectral-Spatial Unified Strategy, particularly its spatial unification component, which effectively processes varying spatial resolutions, and the Local-Global Window Attention mechanism. The LGWA module, with its ability to capture both fine-grained local details and broader contextual information, is particularly adept at delineating intricate building boundaries and accurately segmenting buildings of diverse sizes.
@@ IV-E1 Building
TSSUN 95.71 94.87 95.29 91.00
For the bi-temporal semantic change detection task on the WHU-CD dataset, TSSUN demonstrates leading performance against other SOTA methods, as shown in Table IV. The WHU-CD dataset demands accurate identification of ’from-to’ semantic transitions between two time points. TSSUN excels here due to its inherent design for unified temporal and spectral modeling. The TSSUS component allows the network to effectively learn temporal evolutionary patterns and relationships between different land cover classes, while the LGWA mechanism enhances the discrimination of subtle yet significant semantic changes from unchanged areas, leading to more precise change maps.
@@
RSM-CD 93.07 91.20 92.13 85.40
In the context of bi-temporal binary change detection on the LEVIR-CD dataset, which features a large number of buildings of various sizes and styles undergoing changes, TSSUN surpasses existing methods. Table V shows that our method achieves the highest F1-score of 91.59% and an IoU of 84.49%. The strength of TSSUN on this dataset lies in its robust temporal modeling capabilities, facilitated by TSSUS, which allows for consistent feature representation across different time points. Furthermore, the LGWA mechanism’s proficiency in extracting salient local changes while considering global context ensures high accuracy in detecting both small and large-scale building changes, minimizing missed detections and false alarms often encountered with heterogeneous scene elements.
@@
TSSUN 93.17 90.07 91.59 84.49
Finally, on the TSCD dataset, which presents a challenging multi-temporal binary change detection scenario with longer image sequences, TSSUN achieves the best results, as indicated in Table VI, with an F1-score of 66.48% and an IoU of 49.79%. The TSCD dataset requires robust modeling of temporal dependencies across multiple observations. TSSUS is specifically designed to handle inputs with varying temporal lengths and to model the continuous evolution of land cover. This, combined with the LGWA’s capacity to aggregate contextual information effectively over extended temporal sequences, enables TSSUN to accurately identify changes in complex, evolving landscapes where other methods might falter due to the increased temporal dimensionality.
@@
TSSUN 66.48 49.79 98.05
Figure 4 presents a qualitative comparison of our method against other SOTA approaches on representative samples from all four datasets. Visually, TSSUN consistently produces more accurate and complete segmentation and change maps. The results exhibit fewer false positives and false negatives, particularly in challenging areas such as those with small objects, intricate boundaries, or subtle temporal variations. For instance, in building segmentation on the WHU dataset (Figure 4(a)), TSSUN generates sharper edges and more complete building shapes. Similarly, for change detection tasks on WHU-CD (Figure 4(b)), LEVIR-CD (Figure 4(c)), and TSCD (Figure 4(d)), our method demonstrates superior capability in precisely localizing changed regions while maintaining the integrity of unchanged areas. This visual superiority can be attributed to the model’s enhanced capability, derived from the synergistic operation of TSSUS and LGWA, to learn highly discriminative features and effectively model contextual relationships across the temporal, spectral, and spatial dimensions, resulting in outputs that are more coherent and closely aligned with ground truth.
IV-E2 LULC
The efficacy of the proposed Temporal-Spectral-Spatial Unified Network is rigorously evaluated against SOTA methods on two challenging benchmark datasets: LoveDA for single-temporal semantic segmentation and DynamicEarthNet for multi-temporal land cover classification.
On the LoveDA dataset, TSSUN demonstrates superior performance, achieving the highest OA of 71.82% and mIoU of 65.73% as shown in Table VII. This leading performance can be attributed to TSSUN’s architecture, particularly the Local-Global Window Attention mechanism, which effectively captures both fine-grained local details and broader contextual information. This capability is crucial for accurately segmenting multiscale objects and navigating the complex scenes and backgrounds characteristic of the high-resolution LoveDA imagery. Furthermore, the Temporal-Spectral-Spatial Unified Strategy enables robust feature representation by adeptly unifying spatial and spectral information, which is vital for handling the heterogeneity present in such diverse urban and rural landscapes with uneven class distributions.
For the multi-temporal DynamicEarthNet dataset, TSSUN again surpasses existing methods, yielding the top scores across all reported metrics: Semantic Change Segmentation (SCS) score of 29.9, class-agnostic Binary Change score (BC) of 38.9, and mIoU of 54.7 as shown in Table VIII. The success of TSSUN on this dataset underscores the effectiveness of the TSSUS in explicitly modeling the temporal dimension. By leveraging temporal meta-information to ensure consistency across sequences of varying lengths, TSSUN adeptly manages long-sequence image time series and discerns subtle land cover changes despite spectral variations across different time points—a key challenge in DynamicEarthNet. This robust temporal modeling, distinct from its spatial feature extraction strengths highlighted in the LoveDA analysis, allows for consistent and accurate classification and change analysis across the 24 annotated time steps.
@@ IV-E2 LULC
Methods F1-score per category (%) AF OA mIoU
@@
TSSUN 68.72 76.13 78.83 90.21 54.73 67.03 82.67 74.81 71.82 65.73
@@
TSSUN 29.9 38.9 21.8 54.7
@@ V DISCUSSION
The proliferation of diverse remote sensing data sources necessitates models capable of unified processing across varied data characteristics and task requirements. This research confronts the critical challenge of existing deep learning models in remote sensing dense prediction, which are typically constrained to fixed input-output configurations and struggle to generalize across the inherent heterogeneity in temporal length, spectral channels, and spatial resolution of imagery, as well as diverse task specifications such as semantic segmentation, semantic change detection, and binary change detection. This rigidity imposes significant burdens in terms of retraining and fine-tuning for each specific scenario, limiting practical applicability in dynamic operational environments.
To surmount this limitation, we proposed the Temporal-Spectral-Spatial Unified Network, a novel architecture engineered for comprehensive unification. The cornerstone of TSSUN is the Temporal-Spectral-Spatial Unified Strategy, which systematically decouples and standardizes data representations. On the input side, TSSUS leverages dimension-specific meta-information to harmonize temporal sequences of varying lengths, align spectral inputs with differing channel counts, and reconcile features from disparate spatial resolutions. This is achieved by modeling the continuity in temporal evolution, the inherent relationships between spectral bands, and the local similarity in spatial regions. Concurrently, on the output side, TSSUS exploits shared structural elements among the three primary dense prediction tasks. It supports variable temporal output lengths by identifying common temporal patterns, accommodates an arbitrary number of output classes by capturing inter-class relationships, and ensures spatial consistency across resolutions. Furthermore, the integration of a Local-Global Window Attention mechanism enhances TSSUN’s capacity. LGWA efficiently processes local contextual information through overlapping windowed attention blocks and aggregates global context via a subsequent global attention layer, striking a balance between computational load and feature representation power essential for complex remote sensing scenes.
The efficacy of the TSSUN framework in addressing the identified research problem is substantiated by extensive experimental validation. Results across multiple datasets, encompassing diverse Temporal-Spectral-Spatial configurations and tasks like building detection and land use/land cover classification, demonstrate that a single, pre-trained TSSUN model can successfully adapt to heterogeneous inputs and unify these distinct dense prediction tasks, consistently achieving performance comparable to or exceeding state-of-the-art specialized models.
Despite these advancements, the current TSSUN framework possesses certain limitations. Firstly, the TSSUS component relies on the availability and accuracy of explicit meta-information (e.g., acquisition dates, sensor band specifications, ground sampling distance) to guide the unification across temporal, spectral, and spatial dimensions. The performance may be suboptimal if such meta-information is imprecise, incomplete, or unavailable, potentially hindering the model’s adaptability in data-scarce or poorly documented scenarios. Secondly, while the LGWA mechanism is designed for efficiency, the computational demand of its global attention component, although optimized, might still present a scalability challenge when processing extremely large-scale images or exceptionally long and dense temporal sequences, potentially requiring substantial memory and processing resources.
Future research will focus on mitigating these limitations and expanding the capabilities of TSSUN. One promising direction involves developing methods for the implicit learning or unsupervised inference of the requisite meta-information directly from the remote sensing data, thereby reducing dependency on external inputs and enhancing robustness. Another avenue for exploration is the refinement of the LGWA module, possibly by incorporating more sophisticated hierarchical attention structures or adaptive computational pruning techniques to further improve its efficiency and scalability for massive datasets. Additionally, future work could extend the unification paradigm of TSSUN to encompass a broader array of remote sensing tasks and potentially integrate other data modalities, such as SAR or LiDAR, within a unified analytical framework.

### v2

@@ IV-A Datasets
We offer a brief description of the experimental building and LULC scenario datasets in Table I.
1) Building scenario datasets: The WHU Building dataset [37] is divided into two main components: one containing satellite imagery and another composed of aerial photos. In our study, we utilize the aerial photo subset, which consists of 8,189 images. These images are split into 4,736 for training, 1,036 for validation, and 2,416 for testing, each with a spatial resolution of 0.3 meters. In total, this subset represents over 22,000 buildings covering an area in excess of 450 square kilometers. Our experiments were conducted using the original partitioning scheme and image dimensions (512×512) as specified by the WHU dataset.
The WHU-CD dataset [37] includes bitemporal very high-resolution (VHR) aerial images taken in 2012 and 2016, which clearly highlights major changes in building structures. The dataset is partitioned into non-overlapping patches of 1024×1024 pixels. These patches are further allocated into training, validation, and test sets following a 7:1:2 ratio.
@@
2) LULC scenario datasets: The LoveDA dataset [40] consists of 5,987 high-resolution optical remote sensing images (with a ground sampling distance of 0.3 m) each sized at 1024×1024 pixels. It covers seven land cover classes: building, road, water, barren, forest, agriculture, and background. The dataset is divided into 2,522 training images, 1,669 images for validation, and 1,796 images for testing, all drawn from two distinct scenes—urban and rural—from three Chinese cities: Nanjing, Changzhou, and Wuhan. The dataset poses considerable challenges due to the presence of multiscale objects, complex backgrounds, and uneven class distribution.
@@ IV-B Baseline
To evaluate the effectiveness of the proposed STSUN, we conducted comparative experiments with various benchmark methods on building and LULC scene datasets. Since STSUN is adaptable to multiple datasets, it was trained and validated across all datasets in various scenarios, whereas the benchmark methods were trained and validated on individual datasets.
On the four building scene datasets, the compared CNN-based models include FCN [42], SegNet [43], U-Net [44], PSPNet [45], HRNet [46], MA-FCN [47], Deeplabv3+ [48], ResUNet-a [49], MAPNet [50], D-LinkNet [51], SIINet [52], FC-EF [10], FC-Siam-Diff [10], FC-Siam-Conc [10], STANet [38], DTCDSCN [53], SNUNet [54], CDNet [55], DDCNN [56], DASNet [57], DSIFN [58], HANet [59], USSFCNet [60], and SEIFNet [61], the compared transformer-based models include Segformer [62], ChangeFormer [63] and A2Net [64], and the CNN-transformer hybrid models include BDTNet [65], TransUNet [66], CMTFNet [67], BIT [15], MTCNet [68], MSCANet [69], AMTNet-50 [70], Contrast-COUD [39] and TS-COUD [39].
@@ IV-C Implementation Details
2) Training and Inference: The STSUN model was implemented using PyTorch [86] and executed on a single RTX A100 GPU (80G). Due to the heterogeneous image resolutions across the datasets, the batch size was set to 16 for the four building scene datasets and 4 for the two LULC scene datasets. Our optimization strategy combined binary cross-entropy loss with Dice coefficient loss, facilitating a balanced performance optimization. The AdamW optimizer [87] was initialized with a learning rate of 0.0001 and a weight decay of 0.001. A learning rate scheduler was employed to reduce the learning rate by a factor of 0.1 if no increase in the mean F1-score was observed on the aggregate validation set for 5 consecutive epochs. The training process spanned 100 epochs, ensuring robust convergence, and the best performing checkpoints—corresponding to the maximum mean F1-scores achieved—were retained for the testing phase. Furthermore, in order to ensure comparability with existing methodologies, all models were initialized using the default PyTorch settings across all datasets.
@@ IV-D Ablation Study
To ascertain the efficacy of the proposed unified strategy of STSUN and LGWA, ablation studies were performed on the TSCD dataset. The unified strategy of STSUN is denoted as ”decoupled unification,” which consistently preserves the independence of the temporal dimension throughout the unification process across various dimensions. As comparative baselines, the decoupled unification strategy was replaced with a strategy involving the direct unification of spatial, temporal, and spectral dimensions at both the input and output stages (referred to as ”coupled unification”). Specifically, in the input stage, this coupled unification strategy merges the temporal dimension with the channel dimension, without preserving an independent temporal dimension. The DUM is then used to map the combined channel dimension into a unified feature. In the output stage, the temporal dimension is separated from the channel dimension, and the DUM is employed to map the unified feature to the desired temporal length required by the task. Concurrently, LGWA was substituted with standard global attention. This experimental design serves to highlight the distinct advantages of STSUN in its independent extraction of temporal features and the capability of LGWA in the simultaneous capture of both local and global contextual information.
TABLE II: Ablation study on the TSCD Dataset. The Best Values Are Highlighted in Bold.
Strategy Attention F1(%) IoU(%) OA(%)
Coupled Unification Global Attention 62.87 45.85 94.38
Coupled Unification LGWA 63.22 46.22 94.72
Decoupled Unification Global Attention 65.91 49.15 97.21
Decoupled Unification LGWA 66.48 49.79 98.05
Table II displays the results of our ablation studies, demonstrating that both the decoupled unification strategy and LGWA outperform their respective baseline alternatives. Furthermore, their combined use leads to even greater performance improvements.
Specifically, the decoupled unification strategy enhances performance by preserving an independent temporal dimension. This allows STSUN to extract features from remote sensing images at various time steps during the encoder stage. Subsequently, in the feature fusion stage, feature-level fusion effectively reduces interference from irrelevant information [13]. Moreover, when performing specific tasks, this strategy enables more effective integration with task embeddings and allows for the adjustment of temporal length to suit task requirements. This capability facilitates the joint modeling of multiple dense prediction tasks, leveraging data from various tasks to collectively improve their individual performance.
@@ IV-E Overall Comparison
IV-E1 Building Scenario
The efficacy of the proposed Spatial-Temporal-Spectral Unified Network was rigorously evaluated through extensive experiments on four benchmark remote sensing datasets: the WHU dataset for single-temporal semantic segmentation, the WHU-CD dataset for bi-temporal semantic change detection, the LEVIR-CD dataset for bi-temporal binary change detection, and the TSCD dataset for multi-temporal binary change detection. STSUN was compared against several state-of-the-art approaches, with quantitative results summarized in Table III, IV, V and VI.
On the WHU dataset, characterized by its complex building footprints and significant variations in object scale, our proposed STSUN achieves state-of-the-art performance. As detailed in Table III, STSUN obtains the highest IoU of 91.00% and an F1-score of 95.29%. This superior performance can be attributed to STSUN, particularly its spatial unification component, which effectively processes varying image size, and the Local-Global Window Attention mechanism. The LGWA module, with its ability to capture both fine-grained local details and broader contextual information, is particularly adept at delineating intricate building boundaries and accurately segmenting buildings of diverse sizes.
@@ IV-E1 Building Scenario
STSUN 95.71 94.87 95.29 91.00
For the bi-temporal semantic change detection task on the WHU-CD dataset, STSUN demonstrates leading performance against other SOTA methods, as shown in Table IV. The WHU-CD dataset demands accurate identification of ’from-to’ semantic transitions between two time points. STSUN excels here due to its inherent design for unified temporal and spectral modeling. The STSUN component allows the network to effectively learn temporal evolutionary patterns and relationships between buildings and other objects, leading to more precise change maps.
@@
STSUN 93.07 91.20 92.13 85.40
In the context of bi-temporal binary change detection on the LEVIR-CD dataset, which features a large number of buildings of various sizes and styles undergoing changes, STSUN surpasses compared methods. Table V shows that STSUN achieves the highest F1-score of 91.59% and an IoU of 84.49%. The strength of STSUN on this dataset lies in its robust temporal modeling capabilities, which allows for consistent feature representation across different time points. Furthermore, the LGWA mechanism’s proficiency in extracting salient local changes while considering global context ensures high accuracy in detecting both small and large-scale building changes, minimizing missed detections and false alarms often encountered with heterogeneous scene elements.
@@
STSUN 93.17 90.07 91.59 84.49
Finally, on the TSCD dataset, which presents a challenging multi-temporal binary change detection scenario with longer image sequences, STSUN achieves the best results, as indicated in Table VI, with an F1-score of 66.48% and an IoU of 49.79%. The TSCD dataset requires robust modeling of temporal dependencies across multiple observations. STSUN is specifically designed to handle inputs with varying temporal lengths and to model the continuous evolution of buildings, which enables STSUN to accurately identify changes in complex, evolving landscapes.
@@
STSUN 66.48 49.79 98.05
Figure 5 presents inference results of STSUN from all four datasets. Visually, STSUN consistently produces accurate and complete segmentation and change maps. The results exhibit fewer false positives and false negatives, particularly in challenging areas such as those with small objects, intricate boundaries, or subtle temporal variations. For instance, in building segmentation on the WHU dataset (Figure 5(a)), STSUN generates sharp edges and complete building shapes. Similarly, for change detection tasks on WHU-CD (Figure 5(b)), LEVIR-CD (Figure 5(c)), and TSCD (Figure 5(d)), STSUN demonstrates superior capability in precisely localizing changed regions while maintaining the integrity of unchanged areas. This visual superiority can be attributed to the model’s enhanced capability, derived from the synergistic operation of STSUN and LGWA, to learn highly discriminative features and effectively model contextual relationships across the spatial, temporal and spectral dimensions, resulting in outputs that are more coherent and closely aligned with ground truth.
IV-E2 LULC Scenario
The efficacy of the proposed Spatial-Temporal-Spectral Unified Network is rigorously evaluated against SOTA methods on two challenging LULC datasets: LoveDA for single-temporal land cover classification and DynamicEarthNet for multi-temporal land cover classification.
On the LoveDA dataset, STSUN demonstrates superior performance, achieving the highest OA of 71.82% and mIoU of 65.73% as shown in Table VII. This leading performance can be attributed to STSUN’s architecture, particularly the Local-Global Window Attention mechanism, which effectively captures both fine-grained local details and broader contextual information. This capability is crucial for accurately segmenting multiscale objects and navigating the complex scenes and backgrounds characteristic of the high-resolution LoveDA imagery.
@@ IV-E2 LULC Scenario
Methods F1-score per category (%) AF(%) OA(%) mIoU(%)
@@
STSUN 68.72 76.13 78.83 90.21 54.73 67.03 82.67 74.81 71.82 65.73
For the multi-temporal DynamicEarthNet dataset, STSUN again surpasses existing methods, yielding the top scores across all reported metrics: SCS score of 29.9, BC of 38.9, and mIoU of 54.7 as shown in Table VIII. The success of STSUN on this dataset underscores the effectiveness of the STSUN in explicitly modeling the temporal dimension. By leveraging temporal metadata to ensure consistency across sequences of varying lengths, STSUN adeptly manages long-sequence image time series and discerns subtle land cover changes despite spectral variations across different time points—a key challenge in DynamicEarthNet.
@@
STSUN 29.9 38.9 21.8 55.7
Figure 6 presents the LULC semantic segmentation results of our proposed STSUN on the single-temporal LoveDA dataset (Figure 6 a) and the multi-temporal Dynamic EarthNet dataset (Figure 6 b). The method achieves strong performance across both, effectively navigating challenges such as the complex spectral-spatial features, multi-scale objects in LoveDA, and the inherent temporal variations within Dynamic EarthNet. This proficiency stems from STSUN’s ability to perform unified representation and modeling of remote sensing data across spatial, temporal and spectral dimensions. Specifically, the STSUN enables harmonized encoding and feature fusion across these dimensions, while the LGWA mechanism efficiently captures both local details and global contextual information, crucial for accurate LULC delineation in diverse scenarios.
Figure 6: Sample inference results on for LULC scene datasets. The input images, ground truths and predictions are shown in the first, second and third rows, respectively. (a) LoveDA dataset sample. (b) DynamicEarthNet dataset sample.
@@ V DISCUSSION
V-A Unified Multitask Learning
STSUN introduces a unified framework for dense prediction, capable of concurrently addressing semantic segmentation, binary change detection, and semantic change detection within a single model. This unification is principally achieved through the trainable task embeddings and the Dimension Unification Module. This approach aligns with the core tenets of multi-task learning (MTL), which posit that jointly learning related tasks can lead to improved generalization by leveraging shared representations [90]. By training on a comprehensive dataset aggregated from these distinct tasks, our model develops more robust features than models trained in a single-task paradigm [91].
Moreover, our framework marks a departure from conventional MTL applications in the remote sensing domain. Prevailing methods typically depend on the availability of paired datasets, where each geographical location has labels for all tasks, and commonly employ multiple task-specific decoder heads to generate predictions. In stark contrast, the key strengths of our model are its capacity to be trained on non-paired data from disparate tasks and its reliance on a single, shared decoder head for all predictions.
The implications of this design are twofold. First, the ability to utilize non-paired data dramatically expands the pool of usable training imagery, addressing the data scarcity issue in the field. Second, the single shared decoder enforces the learning of a more cohesive and generalized feature representation, promoting more effective knowledge transfer between tasks. This results in a framework that is not only more data-efficient but also exhibits enhanced adaptability and robustness across a diverse range of dense prediction challenges.
V-B Limitation and Expectation
The increasing heterogeneity of remote sensing data across spatial, temporal, and spectral dimensions presents a fundamental challenge to conventional deep learning models, which are typically engineered for fixed data configurations and singular tasks. This rigidity impedes the integrated analysis of multi-source data, such as fusing Landsat and Sentinel imagery for longitudinal land cover analysis or adapting a single model for both routine monitoring and rapid disaster assessment. Our work addresses this critical gap by introducing a unified framework, STSUN, which demonstrates that a single model can adeptly navigate these STS variations. The scientific significance of this approach lies in its ability to holistically leverage the vast and disparate archives of remote sensing data, fostering the development of more robust, flexible, and data-efficient solutions for dense prediction.
Our experimental results robustly demonstrate that a single STSUN can successfully adapt to heterogeneous inputs and unify these distinct dense prediction tasks. Across multiple datasets, STSUN consistently achieved performance comparable to or exceeding that of state-of-the-art specialized models, highlighting its significant potential for streamlining complex remote sensing data analysis workflows and reducing the need for task-specific model development. The superior performance of STSUN, as evidenced by our extensive experiments, stems from its departure from traditional multi-task learning paradigms. Prevailing MTL methods in remote sensing often require co-registered, fully labeled datasets for all tasks and rely on multiple task-specific decoders. In contrast, STSUN’s architecture, centered on a shared decoder modulated by task and category embeddings, facilitates joint modeling on a diverse corpus of non-paired data aggregated from semantic segmentation, binary change detection, and semantic change detection. Training on this wider spectrum of data forces the model to learn a more generalized and powerful feature representation by extracting shared knowledge across tasks and data sources, leading to enhanced performance without the need for task-specific architectural components.
Despite its promising results, the current implementation of STSUN has two principal limitations that suggest clear avenues for future research. First, the framework’s flexibility is contingent upon explicit metadata, such as predefined task and category embeddings, which must be provided at inference time. A compelling future direction involves exploring methods for the model to implicitly infer the task and desired output configuration directly from the input data context, thereby increasing its autonomy. Second, the potential of our unified modeling approach has been demonstrated only within a supervised learning context. A significant opportunity exists to extend this framework to the self-supervised domain to construct vision foundation models and vision language models. By pre-training the vision model on massive, unlabeled remote sensing archives with diverse STS characteristics, one could develop a powerful remote sensing vision foundation model. Such a model, pre-trained to understand the fundamental structure of STS data, could then be fine-tuned with minimal data for a wide array of downstream applications or potentially used in zero-shot settings, drastically reducing the dependence on labeled data and advancing the capabilities of large-scale Earth observation. At the same time, applying this method to remote sensing vision-language models could enable the models to be trained with large amounts of remote sensing image-text pair data without needing to consider the dimensionality settings of the remote sensing images. This would not only enhance the performance of remote sensing vision-language models but also make them adaptable to various STS settings for remote sensing image and text pair data.